OpenStudy (anonymous):

Graph the function. y = -x + x2

OpenStudy (anonymous):

is that x squared or 2x?

OpenStudy (anonymous):

x squared

OpenStudy (ash2326):

@kaye<33 do you know parabolas?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

y=x2-x=x2-x +1/4-1/4=(x-1/2)^2 -1/4

OpenStudy (anonymous):